It has been known that a fixed dose of a certain drug results to an average increase of pulse rate by at least 12 beats per minute with a standard deviation of 5. A group of 20 patients given the same dose showed the following increases: 15, 12, 10, 8, 14, 15, 16, 11, 7, 13, 9, 10, 12, 11, 9, 10, 17, 14, 15, 7. Is there proof to show that this group has a lower average increase of pulse rate than the ones in general? Use 0.05 level of significance.

I.                Formulate hypothesis: (20pts)

H0:

H1:

 

II.            What is the given level of significance? (5pts)


III.        What is the appropriate test statistics (formula)? (10pts)


IV.        Write the given and your solution (kindly write the CV and TV) 



1
Expert's answer
2021-11-24T17:16:21-0500
"15, 12, 10, 8, 14, 15, 16, 11, 7, 13,""9, 10, 12, 11, 9, 10, 17, 14, 15, 7.""\bar{x}=\dfrac{1}{n}\sum _i x_i=\dfrac{235}{20}=11.75""s^2=\dfrac{1}{n-1}\sum _i (x_i-\bar{x})^2=\dfrac{173.75}{19}""s=\sqrt{s}=\sqrt{\dfrac{173.75}{19}}\approx3.0240266"

I. The following null and alternative hypotheses need to be tested:

"H_0:\mu\geq12"

"H_1:\mu<12"

This corresponds to a left-tailed test, for which a z-test for one mean, with known population standard deviation will be used.


II. Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a left-tailed test is "z_c = -1.6449."

The rejection region for this left-tailed test is "R = \{z: z < -1.6449\}."


III. The z-statistic is computed as follows:



"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{11.75-12}{5/\sqrt{20}}\approx-0.2236"

IV. The 95% confidence interval is



"CI=(11.75-1.96\times\dfrac{5}{\sqrt{20}},11.75+1.96\times\dfrac{5}{\sqrt{20}})""=(9.559, 13.941)"


Critical value is "z_c=-1.6449"

Test value is "z=-0.2236"

Since it is observed that "z = -0.2236 \geq-1.6449=z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach: The p-value is "p=P(Z<-0.2236)=0.411534," and since "p=0.411534>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu" is less than "12," at the "\alpha = 0.05" significance level.

Therefore, there is not enough evidence to claim that this group has a lower average increase of pulse rate than the ones in general, at the "\alpha = 0.05" significance level.